OpenStudy (jmansedg):

Using the Pythagorean theorem, find the measure of the hypotenuse when a = 3.3cm and b = 2.8cm. A. 4.3cm B. 6.1cm C. 7.2cm D. 8.0cm

OpenStudy (foreverandalways13):

do you know the theorem?

OpenStudy (foreverandalways13):

and how it works?

OpenStudy (jmansedg):

nope

OpenStudy (jmansedg):

@silveralchemist09

OpenStudy (cloverracer):

\[a^2+b^2=c^2\] is the pythagoras theorem. Plug in and solve.

OpenStudy (foreverandalways13):

okay so the theorem is a^2+b^2=c^2 and it only works for right triangles. so you have to plug in the numbers to the correct variable for example if a=2 and b=4 then c=2 root 5 so try it for this problem

OpenStudy (cloverracer):

\[3.3cm^2+2.8cm^2=c^2\]

OpenStudy (foreverandalways13):

yep and this is where a calculator becomes helpful :)

OpenStudy (cloverracer):

We showed you what is needed to be plugged in, now find out what 3.3^2 equals and what 2.8cm^2 equals.

OpenStudy (jmansedg):

b?

OpenStudy (cloverracer):

after you found those, you should get.. \[10.89 + 7.84 = c^2\]

OpenStudy (foreverandalways13):

exactly now solve for c

OpenStudy (cloverracer):

\[10.89 + 7.84 = 18.73\] now square 18.73

OpenStudy (foreverandalways13):

yep and the answer should be 2.3

OpenStudy (foreverandalways13):

does that make sense or do u need another example?

OpenStudy (cloverracer):

The square root of 18.73 is rounded to 4.3.

OpenStudy (foreverandalways13):

oops sorry typo

OpenStudy (cloverracer):

which would be A.

OpenStudy (foreverandalways13):

yep!

OpenStudy (cloverracer):

@jmansedg you should learn how to plug in and solve on your own as you will be using this for the rest of your math career. believe me.

OpenStudy (jmansedg):

If marc TMH = 268°, find the measure of the corresponding minor arc. circle Z with radius ZH, diameter MT, and point A opposite point H A. 90° B. 268° C. 134° D. 92°

OpenStudy (foreverandalways13):

@CloverRacer ur completely right...im in precalc and still use the pythagorean theorem

OpenStudy (cloverracer):

Please do not post questions from your Geometry exam @jmansedg

OpenStudy (jmansedg):

i am not

OpenStudy (cloverracer):

@foreverandalways13 I like your username, i'm a swiftie as well :) I am in algebra II and if you forget the slightest thing, your whole 3 page equation is wrong.

OpenStudy (cloverracer):

In the URL, it says geometry 7 exam..

OpenStudy (foreverandalways13):

yeah no kidding! and yay a fellow swiftie!!!

OpenStudy (jmansedg):

it does that trust me is it not

OpenStudy (foreverandalways13):

sorry but if we help u we might get in trouble because thats what it looks like

OpenStudy (jmansedg):

it is not trust me please

OpenStudy (cloverracer):

It literally says 'exam' twice in the URL. I'm sure your teacher has gone over this in a teachlet as you attend roads.advancedmathematics.

OpenStudy (foreverandalways13):

its kinda a risk i dont want to take...sorry :(

OpenStudy (jmansedg):

ok